# RouteKG: A knowledge graph-based framework for route prediction on road networks

### v3

@@ Abstract
Short-term route prediction on road networks allows us to anticipate the future trajectories of road users, enabling various applications ranging from dynamic traffic control to personalized navigation. Despite recent advances in this area, existing methods focus primarily on learning sequential transition patterns, neglecting the inherent spatial relations in road networks that can affect human routing decisions. To fill this gap, this paper introduces RouteKG, a novel Knowledge Graph-based framework for route prediction. Specifically, we construct a Knowledge Graph on the road network to encode spatial relations, especially moving directions that are crucial for human navigation. Moreover, an n-ary tree-based algorithm is introduced to efficiently generate top-K routes in a batch mode, enhancing computational efficiency. To further optimize the prediction performance, a rank refinement module is incorporated to fine-tune the candidate route rankings. The model performance is evaluated using two real-world vehicle trajectory datasets from two Chinese cities under various practical scenarios. The results demonstrate a significant improvement in accuracy over baseline methods. We further validate the proposed method by utilizing the pre-trained model as a simulator for real-time traffic flow estimation at the link level. RouteKG holds great potential for transforming vehicle navigation, traffic management, and a variety of intelligent transportation tasks, playing a crucial role in advancing the core foundation of intelligent and connected urban systems.
Index Terms:
Route prediction, Road network representation, Knowledge graph, Intelligent transportation systems
@@ I Introduction
In intelligent transportation and urban systems, with the increasing prevalence of mobile sensors (e.g., GPS devices) and vehicular communication technologies, the ability to predict road users’ future routes is not merely a convenience but a necessity to support a range of applications such as vehicle navigation [1], traffic management [2], accident prediction [3] and location-based recommendation [4, 5, 6]. There are generally two types of route prediction tasks. On the one hand, for transport planning applications, it is often required to predict the complete route (as a sequence of road links) from the origin to the destination. This is typically referred to as route choice modeling in the literature [7], where the destination information has to be given. On the other hand, for real-time ITS applications, the destination information may not be available, and it is usually adequate to predict the near-future route trajectory of a moving agent based on the observed trajectory so far. This study focuses on the latter, which we call the short-term route prediction problem.
@@
Numerous methods have been proposed in the literature to tackle this problem. Recent works typically use Recurrent Neural Networks (RNNs) [8], especially Long Short-Term Memory (LSTM) [9] and Gated Recurrent Unit (GRU) [10], to capture sequential dependencies in trajectory data [11, 12]. Most existing models focus primarily on learning sequential patterns for route prediction, often overlooking the inherent spatial structure in road networks that can affect human routing decisions. To address this issue, some studies have started to leverage Graph Neural Networks (GNNs) [13] to encode road networks for improved prediction of vehicle trajectories [14] and traffic conditions [15, 16, 17]. However, these methods still treat road networks merely as generic graphs, oversimplifying their structure and disregarding crucial geographical attributes and spatial factors.
@@
As a type of spatial network, road networks consist of a set of spatial entities (e.g., intersections, links, etc.) organized in a way to facilitate traffic flows in a mostly 2-dimensional space. The relationships between these entities can be described by a set of spatial factors such as direction, distance, and connectivity. For example, one of the important spatial factors to consider in routing problems is the direction of travel (i.e., goal direction). It has been widely recognized in the navigation and cognitive psychology literature that humans utilize directional cues to navigate their environment [18, 19]. Existing short-term route prediction models, however, often neglect the inherent static spatial structural relationships and the dynamic relations between these networks and human-intended moving directions, thus possibly leading to sub-optimal model performance. These limitations highlight the need for a more spatially explicit model to accurately learn and incorporate spatial relationships and human-intended moving directions within road networks.
@@
There are other challenges for short-term route prediction on road networks. Firstly, most existing methods focus on generating a single predicted route [20, 21]. However, due to the inherent uncertainties, providing multiple route predictions can have more practical implications. For instance, traffic managers can optimize real-time traffic flow by considering multiple potential routes of moving vehicles, and transportation system users can benefit from having a wider variety of routing options. Secondly, as road networks grow in size, scalability becomes a challenge for GNN-based methods [22], as they require substantial computational and memory resources, limiting their applicability to large-scale networks. Lastly, the prediction performance is heavily dependent on the availability of information about the destination (or goal). Generally, potential performance enhancements can be achieved by incorporating goal information into route prediction models. The availability of goal information can vary, including (1) no information, (2) goal direction only, and (3) complete goal information. These varying degrees of goal information availability can impact the route prediction to different extents, but no existing studies have carried out a comprehensive evaluation across all these scenarios.
@@
With the aforementioned challenges, we propose a novel model, “RouteKG”, which leverages the potential of Knowledge Graphs (KGs) [23] to encode road networks for short-term route prediction. Unlike existing models that rely on sequence-to-sequence (seq2seq) structures [24], our approach interprets route prediction as a Knowledge Graph Completion (KGC) task [25]. Specifically, we propose a Knowledge Graph Module that can predict the future links (tail entities) a user might traverse based on current links (head entities) and moving directions (relations) without solely relying on sequential structures. The module explicitly incorporates the goal moving direction (estimated or actual) into the future route prediction process, better aligning with the intrinsic nature of human navigation. In addition, we employ a Route Generation Module to efficiently generate top-$K$ route candidates, and a Rank Refinement Module that can model the dependencies between different links within each predicted route to rerank the route candidates for their consistency, resulting in the final top-$K$ predictions. Our proposed KG-based framework can effectively model the spatial relations, thus outperforming existing baselines by a large margin, and could benefiting a range of other transportation or routing tasks. To summarize, this paper contributes to the literature as follows:
@@
  • We introduce RouteKG, a novel KG-based modeling framework for short-term route prediction. In this approach, we adapt the KG to represent road networks and reformulate the route prediction problem as a KGC task. Then, the learned road network and route representations can enhance prediction accuracy and interpretability.
@@
  • We propose an $n$-ary tree-based route generation algorithm that enables efficient batch generation of future routes based on predicted probabilities derived from the KG. Additionally, we employ a rank refinement module that effectively prioritizes routes for their consistency by modeling dependencies between their road links, resulting in more accurate, trustworthy, and reliable top-$K$ route predictions.
@@
  • Through extensive experiments on two real-world vehicle trajectory datasets from Chengdu and Shanghai, the results demonstrate the superior prediction performance of RouteKG over state-of-the-art baseline models across various scenarios of goal information availability, with low response latency. Furthermore, a case study using the trained RouteKG as a simulator to estimate real-time traffic flows at the link level demonstrates our method’s effectiveness in diverse application scenarios.
@@ II-A1 Motion Prediction
Classic methods leverage mathematical models grounded in physics and geometry to focus on the deterministic aspects of an agent’s motion, offering simplicity, interpretability, and efficiency [30]. Yet, these methods struggle to capture the stochastic behavior of agents in complex environments [31].
On the other hand, deep learning-based motion prediction methods leverage neural networks’ power to model the complexities of agent behavior [11]. These methods aim to learn the intricate, often non-linear, relationships between different influencing factors from large-scale data. Approaches such as RNNs and Generative Adversarial Networks (GANs) are commonly used [32, 33, 34]. Recent efforts employ diffusion process [35] simulate the process of human motion variation from indeterminate to determinate [36]. The advantage of deep learning methods is their ability to capture the underlying patterns and subtleties that traditional mathematical models might miss. However, they require extensive computational resources and large amounts of training data, and often lack the interpretability of classic methods [28].
@@ II-A2 Route Prediction
Route prediction, distinct from motion prediction, forecasts the future trajectories of agents that typically operate within road network constraints, necessitating different problem formulations and solutions. Similar to motion prediction, models designed for short-term route prediction can also be broadly classified into traditional approaches and deep learning-based methods.
Traditional methods utilize shortest path-based methods such as Dijkstra’s algorithm [37], Bellman-Ford, and A* [38] for route prediction tasks. However, these dynamic programming-based methods require destination information to generate potential routes for trajectory prediction. As the destination information is often unavailable for short-term route prediction, other works have employed Kalman Filters [39] or Hidden Markov Models (HMMs) [40, 41] to predict users’ destinations and routes. Nevertheless, these methods struggle to model long-term temporal dependencies due to relatively simple model structures.
@@ II-B1 Knowledge Graph Completion
To summarize, KGC is a process that leverages machine learning to infer and predict missing knowledge automatically. It leverages the rich structure of KGs, employing effective entity and relation representations for improved prediction.
@@ II-B2 Mobility Knowledge Graph
KGs have been increasingly utilized to address complex urban mobility problems. Mobility KGs have witnessed considerable growth and advancements in recent years, particularly with integrating multi-source transportation data, creating KGs derived from GPS trajectory data, and utilizing structured knowledge bases to augment urban mobility data analysis.
[60] devised a KG for urban traffic systems to uncover the implicit relationships amongst traffic entities and thereby unearth valuable traffic knowledge. Similarly, [61] constructed an urban movement KG using GPS trajectory data and affirmed the practicality of their model by predicting the level of user attention directed towards various city locations. [62] put forth a generalized framework for multi-source spatiotemporal data analysis, underpinned by KG embedding, intending to discern the network structure and semantic relationships embedded within multi-source spatiotemporal data. Several studies have focused on building KGs grounded on geographical information and human mobilities for various applications, such as predicting subsequent locations (i.e., Point of Interest recommendation) [63, 64, 65], modeling event streams [66], learning user similarity [67], forecasting destinations [68, 69], and performing epidemic contact tracing [70].
@@ III-A Problem Formulation
A map-matched route $x$ of length $m$ is a sequence of links, $x=\left\{e^{1},e^{2},...,e^{m}\right\},x\in\mathcal{X}$. For every consecutive pair of links $(e^{i},e^{i+1})$, there exists a node $v$ in the graph $\mathbf{G}$ that connects the two edges. The set of all map-matched routes is denoted as $\mathcal{X}$. The $i$-th route can be partitioned into an observed route $x^{o}_{i}=\{e^{j}_{i}\}^{\Gamma}_{j=1}$ with length $\Gamma$ and a future route $x^{f}_{i}=\{e^{j}_{i}\}_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}$ with length $\Gamma^{\prime}$, where $\mathcal{X}^{o}=\{x_{i}^{o}\}_{i=1}^{|\mathcal{X}|}$ and $\mathcal{X}^{f}=\{x_{i}^{f}\}_{i=1}^{|\mathcal{X}|}$.
@@
[$\left[\left\{x^{o}_{i}\right\}_{i=1}^{|\mathcal{X}|};\mathbf{G}\right]\stackrel{{\scriptstyle\mathcal{F}_{1}(\cdot;\Theta_{1})}}{{\longrightarrow}}\{x^{f}_{i}\}_{i=1}^{|\mathcal{X}|},$ (1)
Subproblem 2 (Route prediction with goal direction only $\mathcal{F}_{2}$) The goal direction $r^{d}_{i}$ is known in addition to the observed routes. The mapping function $\mathcal{F}_{2}$ leverages the goal direction to predict the future routes more accurately:
[$\left[\left\{x^{o}_{i};r^{d}_{i}\right\}_{i=1}^{|\mathcal{X}|};\mathbf{G}\right]\stackrel{{\scriptstyle\mathcal{F}_{2}(\cdot;\Theta_{2})}}{{\longrightarrow}}\{x^{f}_{i}\}_{i=1}^{|\mathcal{X}|},$ (2)
@@
[$\left[\left\{x^{o}_{i};r^{d}_{i};e^{\Gamma+\Gamma^{\prime}}_{i}\right\}_{i=1}^{|\mathcal{X}|};\mathbf{G}\right]\stackrel{{\scriptstyle\mathcal{F}_{3}(\cdot;\Theta_{3})}}{{\longrightarrow}}\{x^{f}_{i}\}_{i=1}^{|\mathcal{X}|},$ (3)
@@ III-B Knowledge Graph
A KG is a heterogeneous structured data representation containing entities (nodes) and their interrelations (edges). The edges carry precise semantic information about the relation type or associated attributes. Formally, the graph is often represented by triplets: $\mathcal{G}=\left\{(h,r,t)\text{ }|\text{ }h,t\in\mathcal{E},r\in\mathcal{R}\right\}$, where $h$ represents the head entity, $r$ the relation, and $t$ the tail entity. $\mathcal{E}$ is the set of entities, and $\mathcal{R}$ the set of relations. These triplets concisely encode factual information for efficient knowledge discovery, inference, and integration. The graph not merely serves as a repository of existing knowledge but also facilitates the inference of missing information. This process, known as Knowledge Graph Completion (KGC), finds a tail entity $\hat{t}$ given a head entity and a relation, denoted as $(h,r,\hat{t})$, or its reverse, denoted as $(\hat{h},r,t)$, thereby completing a partial triplet.
@@ IV-A RouteKG Framework Overview
Next, the Knowledge Graph Module is the core component of the proposed model, it takes the observed routes $\mathcal{X}^{o}$ and road network $\mathbf{G}$ to predict future routes. It constructs a knowledge graph on the road network $\mathbf{G}$, learns spatial relations $\mathcal{R}$, and predicts future routes, converting $\mathcal{X}^{o}$ to future route probabilities $\mathrm{Pr}(\widetilde{\mathcal{X}^{f}})$ via $\mathrm{Pr}(\widetilde{\mathcal{X}^{f}}),\mathcal{R}=\mathcal{M}_{kg}\left(\mathcal{X}^{o},\mathbf{G},\mathbf{D};\Theta_{kg}\right)$, where $\Theta_{kg}$ are the module’s parameters, and $\mathcal{R}$ represents the learned spatial relations.
With the future route probabilities $\mathrm{Pr}(\widetilde{\mathcal{X}^{f}})$ in hand, the Route Generation Module employs an $n$-ary tree algorithm to generate potential future routes, yielding the top-$K$ preliminary route predictions. This step is captured by $\{\widetilde{\mathcal{X}^{f}_{k}}\}_{k=1}^{K}=\mathcal{M}_{g}(\mathrm{Pr}(\widetilde{\mathcal{X}^{f}}),\mathbf{G},\mathbf{A})$, where $\widetilde{\mathcal{X}^{f}_{k}}$ denotes the $k$-th generated future route.
When predicting future routes, predicted road links at different time steps are not independent but related. Thus, a Rank Refinement Module is utilized to collectively learn and assess the predicted route. It takes the initial top-$K$ predictions, $\{\widetilde{\mathcal{X}^{f}_{k}}\}_{k=1}^{K}$, and refines them using the spatial relations, $\mathcal{R}$. This refinement is achieved by the mapping $\{\widehat{\mathcal{X}^{f}_{k}}\}_{k=1}^{K}=\mathcal{M}_{r}(\{\widetilde{\mathcal{X}^{f}_{k}}\}_{k=1}^{K},\mathcal{R};\Theta_{r})$, where $\Theta_{r}$ are the module’s parameters. This stage ensures that the final route predictions are accurate by considering the sequence of routes and spatial relations.

### v4

@@ Abstract
Short-term route prediction on road networks allows us to anticipate the future trajectories of road users, enabling various applications ranging from dynamic traffic control to personalized navigation. Despite recent advances in this area, existing methods focus primarily on learning sequential transition patterns, neglecting the inherent spatial relations in road networks that can affect human routing decisions. To fill this gap, this paper introduces RouteKG, a novel Knowledge Graph-based framework for route prediction. Specifically, we construct a Knowledge Graph on the road network to encode spatial relations, especially moving directions that are crucial for human navigation. Moreover, an n-ary tree-based algorithm is introduced to efficiently generate top-K routes in batch mode, enhancing computational efficiency. To further optimize prediction performance, a rank refinement module is incorporated to fine-tune candidate route rankings. The model performance is evaluated using two real-world vehicle trajectory datasets from two Chinese cities under various practical scenarios. The results demonstrate a significant improvement in accuracy over the baseline methods. We further validate the proposed method by utilizing the pre-trained model as a simulator for real-time traffic flow estimation at the link level. RouteKG has great potential to transform vehicle navigation, traffic management, and a variety of intelligent transportation tasks, playing a crucial role in advancing the core foundation of intelligent and connected urban systems. The source codes of RouteKG are available at https://github.com/YihongT/RouteKG.
@@ I Introduction
In intelligent transportation and urban systems, with the increasing prevalence of mobile sensors (e.g., GPS devices) and vehicular communication technologies, the ability to predict road users’ future routes is not merely a convenience but a necessity to support a range of applications such as vehicle navigation [1], traffic management [2], accident prediction [3] and location-based recommendation [4, 5]. There are generally two types of route prediction tasks. On the one hand, for transport planning applications, it is often required to predict the complete route (as a sequence of road links) from the origin to the destination. This is typically referred to as route choice modeling in the literature [6], where information about the destination (or goal) must be given. On the other hand, for real-time ITS applications, goal information may not be available, and it is usually adequate to predict the near-future route trajectory of a moving agent based on the observed trajectory so far. This study focuses on the latter, which we call the short-term route prediction problem.
@@
Numerous methods have been proposed in the literature to address this problem. Recent works typically use Recurrent Neural Networks (RNNs) [7], especially Long Short-Term Memory (LSTM) [8] and Gated Recurrent Unit (GRU) [9], to capture sequential dependencies in trajectory data [10, 11]. Most existing models focus primarily on learning sequential patterns for route prediction, often overlooking the inherent spatial structure of road networks that can affect human routing decisions. To address this issue, some studies have started to use Graph Neural Networks (GNNs) [12] to encode road networks for improved prediction of vehicle trajectories [13] and traffic conditions [14, 15]. However, these methods still treat road networks merely as generic graphs, oversimplifying their structure and disregarding crucial spatial factors.
@@
As a type of spatial network, road networks consist of a set of spatial entities (e.g., intersections, links, etc.) organized in a way to facilitate traffic flows in a mostly 2-dimensional space. The relationships between these entities can be described by a set of spatial factors such as direction, distance, and connectivity. For example, one of the important spatial factors to consider in routing problems is the direction of travel (i.e., goal direction). It has been widely recognized in the navigation and cognitive psychology literature that humans utilize directional cues to navigate their environment [16, 17]. Existing short-term route prediction models, however, often neglect the spatial structural information of road networks and human-intended moving directions, thus possibly leading to sub-optimal model performance. These limitations highlight the need for a more spatially explicit model to incorporate such spatial relationships that influence human routing behavior.
@@
There are other challenges for short-term route prediction on road networks. Firstly, most existing methods focus on generating a single predicted route [18, 19]. However, due to the inherent uncertainties, providing multiple route predictions can have more practical implications. For instance, traffic managers can optimize real-time traffic flow by considering multiple potential routes of moving vehicles, and road users can benefit from having a wider variety of routing options. Secondly, as road networks grow in size, scalability becomes a challenge for GNN-based methods [20], as they require substantial computational and memory resources. Lastly, model prediction performance is heavily dependent on the availability of goal information. Generally, better performance can be achieved by incorporating more goal information into the model. However, the availability of goal information varies, including (1) no information available, (2) only goal direction known, and (3) complete goal information known. These varying degrees of goal information availability can affect route prediction to different extents, but a comprehensive evaluation across all these scenarios is missing.
@@
With the aforementioned challenges, we propose a novel model, RouteKG, which leverages Knowledge Graphs (KGs) [21] to encode road networks for short-term route prediction. Unlike existing models that rely on sequence-to-sequence (seq2seq) structures [22], RouteKG interprets route prediction as a Knowledge Graph Completion (KGC) task [23]. Specifically, we propose a Knowledge Graph Module that can predict the future links (tail entities) a user might traverse based on current links (head entities) and moving directions (relations) without solely relying on seq2seq structures. The module explicitly incorporates the goal moving direction (estimated or actual) into the future route prediction process, better aligning with the intrinsic nature of human navigation. In addition, we employ a Route Generation Module to efficiently generate top-$K$ route candidates, and a Rank Refinement Module that can model the dependencies between different links within each predicted route to rerank the route candidates for consistency, resulting in the final top-$K$ predictions. Our proposed KG-based approach can effectively model the spatial relations, thus outperforming existing baselines by a large margin, and it could benefit a range of transportation or routing tasks. The contributions of this study are summarized as follows:
@@
  • We introduce RouteKG, a novel KG-based modeling framework for short-term route prediction. In this approach, we adapt the KG to represent road networks and reformulate the route prediction problem as a KGC task. Then, the learned road network and route representations can enhance model prediction and interpretability.
@@
  • We propose an $n$-ary tree-based route generation algorithm that enables efficient batch generation of future routes based on predicted probabilities derived from the KG. Additionally, we employ a rank refinement module that effectively prioritizes routes for their consistency by modeling dependencies between road links, resulting in more accurate, trustworthy, and reliable top-$K$ route predictions.
@@
  • Through extensive experiments on two real-world vehicle trajectory datasets from Chengdu and Shanghai, the results validate the superior prediction performance of RouteKG over state-of-the-art baseline models across various scenarios of goal information availability, with low response latency. Furthermore, a case study using the trained RouteKG as a simulator to estimate real-time traffic flows at the link level demonstrates our method’s effectiveness in diverse application scenarios.
@@ II-A1 Motion Prediction
Classic methods utilize mathematical models grounded in physics and geometry to focus on the deterministic aspects of an agent’s motion, offering simplicity, interpretability, and efficiency [28]. Yet, these methods struggle to capture the stochastic behavior of agents in complex environments [29].
On the other hand, deep learning-based motion prediction methods use neural networks to model the complexities of agent behavior [10]. These methods aim to learn the intricate, often non-linear, relationships between different influencing factors from large-scale data. Approaches such as RNNs and Generative Adversarial Networks (GANs) are commonly used [30, 31, 32]. Recent efforts employ diffusion process [33] simulate the process of human motion variation from indeterminate to determinate [34]. The advantage of deep learning methods is their ability to capture the underlying patterns and subtleties that traditional mathematical models might miss. However, they require extensive computational resources and large amounts of training data, and often lack the interpretability of classic methods [26].
@@ II-A2 Route Prediction
Route prediction, distinct from motion prediction, forecasts the future trajectories of agents that typically operate within road network constraints, necessitating different problem formulations and solutions. Similar to motion prediction, models designed for short-term route prediction can also be broadly classified into traditional approaches and deep learning-based methods. Traditional methods utilize shortest path-based methods such as Dijkstra’s algorithm [35], Bellman-Ford, and A* [36] for route prediction tasks. However, these dynamic programming-based methods require destination information to generate potential routes for trajectory prediction. As the destination information is often unavailable for short-term route prediction, other works have employed Kalman Filters [37] or Hidden Markov Models (HMMs) [38, 39] to predict users’ destinations and routes. Nevertheless, these methods struggle to model long-term temporal dependencies due to relatively simple model structures.
@@ II-B1 Knowledge Graph Completion
To summarize, KGC is a process that leverages machine learning to infer and predict missing knowledge automatically. It utilizes the rich structure of KGs, employing effective entity and relation representations for improved prediction.
@@ II-B2 Mobility Knowledge Graph
KGs have been increasingly utilized to address complex urban mobility problems. Mobility KGs have witnessed considerable growth and advancements in recent years, particularly with the integration of multi-source transportation data, creating KGs derived from GPS trajectory data, and utilizing structured knowledge bases to augment urban mobility data analysis. [58] devised a KG for urban traffic systems to uncover the implicit relationships amongst traffic entities and thereby unearth valuable traffic knowledge. Similarly, [59] constructed an urban movement KG using GPS trajectory data and affirmed the practicality of their model by predicting the level of user attention directed towards various city locations. [60] put forth a generalized framework for multi-source spatiotemporal data analysis, underpinned by KG embedding, intending to discern the network structure and semantic relationships embedded within multi-source spatiotemporal data. Several studies have focused on building KGs grounded on geographical information and human mobilities for various applications, such as predicting subsequent locations (i.e., Point of Interest recommendation) [61, 62, 63], modeling event streams [64], learning user similarity [65], forecasting destinations [66, 67], and performing epidemic contact tracing [68].
@@ III-A Problem Formulation
We denote the full set of map-matched routes as $\mathcal{X}=\{x_{i}\}_{i=1}^{|\mathcal{X}|}$. Each route $x\in\mathcal{X}$ is derived from raw GPS trajectories via map-matching [69]. A map-matched route $x$ of length $m$ is thus a sequence of road links, $x=\left\{e^{1},e^{2},...,e^{m}\right\},\ x\in\mathcal{X}$. For every consecutive pair of links $(e^{i},e^{i+1})$, there exists a node $v$ in the road graph $\mathbf{G}$ such that $v$ connects the two edges. The $i$-th route can be partitioned into an observed route $x^{o}_{i}=\{e^{j}_{i}\}^{\Gamma}_{j=1}$ of length $\Gamma$ and a future route $x^{f}_{i}=\{e^{j}_{i}\}_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}$ of length $\Gamma^{\prime}$, where both $\Gamma$ and $\Gamma^{\prime}$ are fixed. The full sets of observed and future routes are denoted as $\mathcal{X}^{o}=\{x_{i}^{o}\}_{i=1}^{|\mathcal{X}|}$ and $\mathcal{X}^{f}=\{x_{i}^{f}\}_{i=1}^{|\mathcal{X}|}$, respectively.
@@
$\left[\left\{x^{o}_{i}\right\}_{i=1}^{|\mathcal{X}|};\mathbf{G}\right]\stackrel{{\scriptstyle\mathcal{F}_{1}(\cdot;\Theta_{1})}}{{\longrightarrow}}\{x^{f}_{i}\}_{i=1}^{|\mathcal{X}|},$ (1)
Subproblem 2 (Route prediction with goal direction only $\mathcal{F}_{2}$) The goal direction $r^{d}_{i}$ (the relative orientation from the last observed road link to the goal link on the road network) is known in addition to the observed routes. The mapping function $\mathcal{F}_{2}$ leverages the goal direction to predict the future routes more accurately:
$\left[\left\{x^{o}_{i};r^{d}_{i}\right\}_{i=1}^{|\mathcal{X}|};\mathbf{G}\right]\stackrel{{\scriptstyle\mathcal{F}_{2}(\cdot;\Theta_{2})}}{{\longrightarrow}}\{x^{f}_{i}\}_{i=1}^{|\mathcal{X}|},$ (2)
@@
$\left[\left\{x^{o}_{i};r^{d}_{i};e^{\Gamma+\Gamma^{\prime}}_{i}\right\}_{i=1}^{|\mathcal{X}|};\mathbf{G}\right]\stackrel{{\scriptstyle\mathcal{F}_{3}(\cdot;\Theta_{3})}}{{\longrightarrow}}\{x^{f}_{i}\}_{i=1}^{|\mathcal{X}|},$ (3)
@@ III-B Knowledge Graph
[…]
@@ IV-A RouteKG Framework Overview
[…]
In future route prediction, predicted road links at different time steps are not independent but related. Thus, the Rank Refinement Module is designed to collectively learn and assess the predicted route. It takes the initial top-$K$ predictions, $\{\widetilde{\mathcal{X}^{f}_{k}}\}_{k=1}^{K}$, and refines them using the spatial relations, $\mathcal{R}$, through the mapping $\{\widehat{\mathcal{X}^{f}_{k}}\}_{k=1}^{K}=\mathcal{M}_{r}(\{\widetilde{\mathcal{X}^{f}_{k}}\}_{k=1}^{K},\mathcal{R};\Theta_{r})$, where $\Theta_{r}$ are the module’s parameters. This stage ensures that the final route predictions are accurate by considering the sequence of routes and spatial relations.